OpenStudy (anonymous):

I NEED HELP WITH 3 QUESTIONS ASAP

OpenStudy (anonymous):

what is it

OpenStudy (anonymous):

are they*

OpenStudy (anonymous):

Solve for h. P = mgh A. h = P – mg B. h=pg/m C.h=p/mg D. h = P + mg

OpenStudy (anonymous):

c

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

just take the variables and put small numbers in

OpenStudy (anonymous):

like 1*2*3=6 right?

OpenStudy (anonymous):

Which literal equations are equivalent to ? Choose all answers that are correct. A. g=w/m B. g=m/w C. w=mg D.w=m/g

OpenStudy (anonymous):

so if 3 is h then 6/1*2 = 3 or h

OpenStudy (anonymous):

equivilent to m=w/g

OpenStudy (anonymous):

c again

OpenStudy (anonymous):

there can be more than one answer btw

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

okay one more

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do you do k12

OpenStudy (anonymous):

Which literal equation is equivalent to 1.4 = kx A. k=x(divided by) 1.4 B. x=1.4(multiplied by) k C.k=1.4(divided by) x D.x=k(divided by) 1.4

OpenStudy (anonymous):

you there

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@StilesStilinski24

OpenStudy (anonymous):

thelast question is above!

OpenStudy (anonymous):

I WILL GIVE MEDAL AND FAN

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

d